Provide Anne with Financial advice on which option has the potential to yield the highest monetary value.Support your rational with calculations using time value of money and comment on the risk return relationship for each option,assume interest rate on savings is 4 %and is compounded semi annually.

Anne is a 55 year old accountant who works at BDO who is about to retire.She has the following decision to make:

Option A -select a lump sum gratuity payment of $120000 with a reduced pension of $1750 per month.

Option B- select a monthly pension of $3300 with no lump sum gratuity payment.

In addition,Anne has a loan $72000 with loan payments of $1200 per month for the next five years.
Answer #1
Savings Interest (Semi Annually) 4%
Lumpsum Gratuity 120000
Loan Outstanding 72000
EMI 1200 pm
Option A Option B Formula
Monthly Pension 1750 3300
Loan Repayment 0 1200
Amount left from monthly receipts 1750 2100
Accummulated balance every 6 months 10500 12600
Amount at the end of 5 years compounded Semiannually 114972 137966 =FV(0.02,10,-10500,0) =FV(0.02,10,-12600,0)
PV of the above discounted at 4% 94499 113398 A =B18/1.04^5 =C18/1.04^5
Lumpsum Gratuity net of Loan repayment 48000 0
Amount at the end of 5 years compounded Semiannually 58512 0 =B20*(1+0.02)^10
PV of the above discounted at 4% 48092.3787 0 B =B21*1/(1+0.04)^5
Total of the PV's of cashflows 142591 113398 A+B
Hence Option A is beneficial. Repaying the loan upright is wise .